Given the following code find the worst case time complexity


binary search (target: integer, a[1..n ]: ascending integers)
k =1
j =n
loop when (k is less than j)
m =floor((k+j)/2)
if (target is larger than the element at m) then k = m+1
else j = m
endloop
if (target equals element at k) then location=k
else location =0
Answer #1
T(n) = T(n/2) + c 
     = T(n/4) + c + c
     = T(n/8) + c + c + c
     ......
     ......
     ......
     = T(n/n) + c + .... + c + c + c [log(n) +1 terms]
     = c + c + .... + c + c + c [log(n) +1 terms]
     = clog(n)
     = Theta(logn)